# Rethinking the Starting Point: Collaborative Pre-Training for Federated Downstream Tasks

### v3

@@ 3.2 CoPreFL in Scenario I (Pre-training with Distributed Clients)
Model-agnostic meta update. Considering the objective function in (4), each participant $j$ downloads the temporary global model $\overline{\Phi^{t}}$ and employs its query set $Q_{j}$ to compute its local query loss $\mathcal{L}_{Q_{j}}(\overline{\Phi^{t}})$, as in line 15 in Algorithm 1. The gradients are also computed locally and sent back to the server, as both are necessary to conduct the meta-update. On the server-side, the overall query meta-loss $\mathcal{L}_{Q}(\overline{\Phi^{t}})$ and the performance variance $\sigma_{Q}^{2}(\overline{\Phi^{t}})$ are computed, according to (3). Then, as described in line 19 of Algorithm 1, CoPreFL updates the temporary pre-training model $\overline{\Phi^{t}}$ through a gradient step with the customized query meta-loss $\mathcal{L}_{meta}$ and the aggregated received gradients, to align it with (4). To derive the meta-loss $\nabla_{\Phi^{t-1}}\mathcal{L}_{meta}(\overline{\Phi^{t}})$, we express it through the chain rule as $\nabla_{\overline{\Phi^{t}}}\mathcal{L}_{meta}(\overline{\Phi^{t}})\times\frac{\partial\overline{\Phi^{t}}}{\partial\Phi^{t-1}}$. Writing $\overline{\Phi^{t}}=\sum_{j\in m}\frac{|S_{j}|}{\sum_{i\in m}|S_{i}|}\Phi^{t,E}_{j}=\sum_{j\in m}\frac{|S_{j}|}{\sum_{i\in m}|S_{i}|}(\Phi^{t,E-1}-\eta\nabla\ell^{E}_{S_{j}}(\Phi^{t}))$, it follows
@@
$\nabla_{\Phi^{t-1}}\mathcal{L}_{meta}(\overline{\Phi^{t}})=\nabla_{\overline{\Phi^{t}}}\mathcal{L}_{meta}(\overline{\Phi^{t}})\times\Big{(}1-\eta\sum_{j\in m}\frac{|S_{j}|}{\sum_{i\in m}|S_{i}|}\frac{\partial}{\partial\Phi^{t-1}}\nabla\ell^{E}_{S_{j}}(\Phi^{t})\Big{)}.$
@@
If we ignore the second derivative term, the meta-loss gradient can be approximated as $\nabla_{\overline{\Phi^{t}}}\mathcal{L}_{meta}(\overline{\Phi^{t}})$. This approach is similar to making a first-order approximation to a meta-update, a common practice in the implementation of MAML variants to reduce complexity [16, 35].
@@
The server then broadcasts the meta-updated global model $\Phi^{t}$ to a new set of participants to begin the next round of pre-training. After $T$ rounds, the final global model $\Phi^{T}$ serves as the pre-trained model for initializing FL in the downstream tasks, i.e., in Figure 1, the set of clients in any downstream task conduct FL starting from the pre-trained model $w^{0}=\Phi^{T}$.
@@
Remark 1 (Key characteristics of CoPreFL meta-update). Using the query datasets, CoPreFL applies a meta-update to the temporary pre-training model, which is a global model obtained through FL on the support datasets. This is a key distinction from existing meta-learning based FL methods discussed in Section 2, which conduct meta updates on client models for personalization. Our method focuses on tailoring this pre-training model to be adaptable to an arbitrary set of downstream FL tasks, whereas existing personalization methods are not concerned with robustness to unseen and heterogeneous data statistics emerging downstream. As we will see in Section 4, this leads to significant improvements of CoPreFL compared with employing these prior methods for pre-training.
@@ 3.3 CoPreFL in Scenario II (Hybrid Client-Server Pre-Training)
We next explore a pre-training scenario where, in addition to participating clients holding data, the server possesses a small dataset $D^{s}$ drawn from the broader population distribution. Unlike in scenario I where we separated participants’ data into support and query sets, viewing the query sets as unseen knowledge to control average performance and balance, in scenario II, we employ all of a client’s samples as support data for local updates. Instead, we treat the server’s data as the query set.
The procedure of CoPreFL for scenario II is detailed in Algorithm 2 in Appendix B. Here, we highlight the key differences from Algorithm 1. First, the temporary pre-training model $\overline{\Phi^{t}}$ is aggregated from local models trained on the each participant’s entire local dataset $D^{p}_{j}$. Second, we facilitate the meta-update of the temporary pre-training model $\overline{\Phi^{t}}$ using the server’s data. In doing so, to help mimic the distributed nature of downstream FL tasks, we randomly partition the dataset $D^{s}$ into $|m|$ equally-sized partitions, from which we can obtain the average loss and variance objectives similar to (3). The temporary global model $\overline{\Phi^{t}}$ is then updated based on meta-loss $\mathcal{L}_{meta}(\overline{\Phi^{t}})$, which is defined similarly to (4) but calculated through meta-updates on the server’s partitioned data.
Note that unequal and/or non-uniformly random partitionings of the server-side dataset for meta-updating could be considered as alternatives, e.g., to force specific data distributions or label splits across partitions. However, due to the server’s lack of prior knowledge about future downstream FL tasks during pre-training, including their dataset sizes and distributions, meta-updating the model with randomly allocated query sets is the most viable solution. We show in Section 4 that this partitioning provides significant performance improvements over other pre-training strategies.
@@
Remark 3. Providing theoretical analysis on the effect of pre-training strategies continues to be an open problem, especially in linking model initialization to downstream task performance. This challenge has existed both in centralized-to-centralized [6, 13, 45, 18, 37, 44] and centralized-to-federated [34, 8] transfers from pre-training to downstream, and persists in the distributed-to-federated case we consider here. We thus leave theoretical analysis of CoPreFL and other strategies to future work, and instead substantiate the effectiveness of our method through extensive experiments. Nevertheless, the effectiveness of CoPreFL can be attributed to the advantage provided by meta-learning, which we employ to obtain robustness across heterogeneous sets of downstream tasks.
@@ 4 Experiments
Table 1: Performance across 10 different non-IID downstream FL tasks, initialized with various non-IID FL pre-training methods. Lowest $X$reports the average accuracy across the clients with the lowest $X$accuracy. CoPreFL provides the best initialization for downstream FL for each scenario, metric, and dataset.
@@ 4.1 Experimental Setup
Datasets and model. For evaluation, consider CIFAR-100 [25], Tiny-ImageNet [26], FEMNIST [5], and PACS [27], adhering to the data splits provided in [38, 35], and adopt ResNet-18 [17]. To model practical scenarios where downstream task labels are unknown during pre-training, for CIFAR-100, the dataset is divided into 80 classes for pre-training and 20 classes for downstream FL tasks, while for Tiny-ImageNet, the dataset is separated into 160 classes for pre-training and 40 classes for downstream FL tasks. Additionally, we explore mixed scenarios where there are overlapping classes between pre-training and downstream tasks. Following [43, 46], we randomly select 95% of the samples from the pre-training dataset for clients, while the remaining 5% form the server dataset. The PACS dataset consists of 4 different data domains, for which we employ a one-domain-leave-out setup [30, 47] between pre-training and downstream FL tasks. More details on the datasets can be found in Appendix C.1. For FEMNIST, we report all results in Appendix D.4.
@@
Pre-training phase. We distribute the pre-training dataset to $|M|=100$ clients following non-IID data partitions according to a Dirichlet distribution [33, 28], and select $|m|=20$ participants out of the $|M|$ clients for each FL round. Results with different $|m|$ and IID setups are reported throughout Appendix D. We adopt a standard approach commonly used in meta-learning-based research [19, 41, 2, 20, 35] for support/query splitting, where we randomly partition each client’s data into 80% support and 20% query sets. See Appendix C.1C.2 for more details.
@@
Downstream FL task and evaluation metrics. To generate each downstream FL task, we randomly select 5 of the 20 classes from the CIFAR-100 dataset and 40 classes from the Tiny-ImageNet dataset, and distribute the corresponding data samples to a set of $|G|=10$ clients following non-IID Dirichlet data distributions (see Appendix D for IID results). Each participant in the downstream phase utilizes 80% of its local data as training samples, while the remaining 20% is reserved for testing samples. We keep the training procedure consistent for each downstream task, using the conventional FedAvg algorithm by default (see Appendix C.2 for detailed settings). For each task, we evaluate the final global model’s performance on test samples from each client $g\in G$, reporting the accuracy and variance of the accuracy distribution across the clients. We consider a total of $X=10$ downstream tasks, and the evaluation metrics are reported as averages (with standard deviations) across the tasks.
@@
Baselines for pre-training. We compare CoPreFL with several established FL algorithms, including (i) standard FedAvg [31], (ii) FedMeta [7], which employs meta-learning for unseen scenarios, and (iii) q-FFL(q > 0) [29], designed to balance performance across clients. When applying these baselines in scenario II, in each pre-training round $t$, after the global model $\Phi^{t}$ has been constructed, we further train with 5 additional iterations on the server dataset. This extended training follows the approach in [43, 3], where the server’s data is used to further refine the global model. Similarly, we introduce a baseline called CoPreFL-SGD, which first constructs a global model according to CoPreFL and then further performs SGD iterations using server data on the global model. Finally, we also consider initializations based on (iv) conventional centralized pre-training [34], and popular FL algorithms like (v) SCAFFOLD [23], (vi) FedDyn [1], and (vii) PerFedAvg [15] (see Table 3(a)).
@@ 4.2 Experimental Results
Main results for scenarios I & II. Tables 1(a) and 1(b) present the test accuracies for CoPreFL across scenarios I and II on the CIFAR-100 and Tiny-ImageNet datasets. In scenario I, CoPreFL demonstrates robust initializations for downstream FL tasks, obtaining a higher average accuracy and reduced variance across clients, and significantly improving accuracies for the worst-performing clients (Lowest 10-20% metrics). This underscores the benefits of balancing the competing objective functions in (3) during meta-updates. See Appendix D.1 for supplemental analyses and results that include varying the number of participants in pre-training, and different data distributions in downstream FL tasks. In scenario II, similarly, CoPreFL consistently outperforms the baselines, in this case by effectively utilizing server data in addition to balancing our objectives (3). The benefit of a small server-side dataset, when available, can be seen through the performance improvements from Table 1(a) to 1(b). Further, the improvement over CoPreFL-SGD suggests that conducting SGD iterations with server data in a centralized manner after meta-updating the global model might divert the pre-trained model away from our designed objectives. This emphasizes the importance of performing meta-learning on the server data following the partitioning outlined in Algorithm 2. For further results on various configurations, including the effects of varying server dataset sizes, see Appendix D.2.
Figure 2: Testing accuracy distributions in different non-IID FL downstream tasks. CoPreFL achieves the best average accuracy (i.e., a distribution leaning toward the right) and smaller performance variance (i.e., a narrower distribution) while also improving the accuracies of worst-performing clients (i.e., on the left tail).
Performance distribution comparison. Figure 2 visualizes more detailed testing accuracy distributions of the final global model on each client in the downstream tasks. We provide visualizations for our method, as well as the methods with the second-best average accuracy and the second-lowest variance from Table 1. CoPreFL’s distributions are narrower in each case, consistent with its reduction in performance variance across clients. Distribution shifts to the right (i.e., higher average accuracy) can also be observed. Importantly, we see that CoPreFL effectively shifts nearly all low-performing clients on the left tail of the baselines towards the right, indicating enhanced accuracy for these clients. Additional results across various scenarios are provided in Appendix D.3.
@@
Effect of balancer $\gamma$ in CoPreFL. Table 2 gives the performance obtained from CoPreFL using different balancers $\gamma$ on Tiny-ImageNet. A larger $\gamma$ implies that the pre-trained model prioritizes the devices’ average performance, whereas a smaller $\gamma$ prioritizes performance balance. In Table 2, as the balancer $\gamma$ increases, we observe an increase in the average accuracy of downstream FL tasks but a higher variance, resulting in performance imbalance. The trend shows that CoPreFL allows control over the relative importance between accuracy and balanced performance within its pre-training environment.
Comparison with other initialization methods. In addition to the baselines in Table 1(b) which consider performance balance or meta-learning, Table 3(a) also evaluates popular algorithms like SCAFFOLD [23], FedDyn [1], and PerFedAvg [15] providing pre-training for scenario I. We see that CoPreFL also outperforms these baselines, further validating our meta-learning approach based on (3). These popular FL algorithms struggle with heterogeneity in unseen downstream tasks and balancing performance. Additionally, we explore initializing downstream FedAvg with random weights as well as a centrally pre-trained model, a concept introduced in [34]. While the centralized method boosts downstream FL accuracy over random initialization, it introduces significant performance variance across clients due to inability to mimic downstream FL characteristics (also seen in Figure 1). More details along with additional results can be found in Appendix D.5.
@@
Centralized [34] 81.30 $\pm$ 2.92 69.44 $\pm$ 2.33
SCAFFOLD [23] 79.15 $\pm$ 3.08 57.84 $\pm$ 1.95
FedDyn [1] 81.23 $\pm$ 2.96 53.17 $\pm$ 2.85
PerFedAvg [15] 81.58 $\pm$ 1.83 49.73 $\pm$ 2.65
@@
Compatibility with other downstream FL algorithms. We next explore the ability of our pre-training method to enhance the performance of downstream FL algorithms other than FedAvg. For this, we consider FedProx [40] and balance-aware q-FFL [29] in each federated downstream task. Table 3(b) shows the results. Overall, we see that CoPreFL consistently achieves superiority in accuracy and variance compared to other pre-training baselines, when combined with different downstream FL algorithms. Details on the implementation and further discussions are provided in Appendix D.6.
@@
Table 4: Results with both seen/unseen classes during downstream FL. CIFAR-100 is adopted in scenario I.
Both unseen/seen classes in downstream FL tasks. In addition to the setting without overlapping classes between pre-training and downstream tasks, we explore a mixed scenario where clients in the downstream FL task also hold “seen classes.” To implement this, we use CIFAR-100 and randomly sampled 10 classes from the pre-training dataset and 10 classes from the downstream dataset, resulting in 10 seen and 10 unseen downstream classes. From this, we constructed 10 downstream tasks by randomly selecting 5 classes in each case, to conduct non-IID downstream FedAvg. The pre-trained models were solely trained on the original 80 classes of CIFAR-100. Table 4 shows the results. We see that the accuracies are generally higher compared to those in Table 1(a), as the downstream tasks involve classes seen during pre-training. Once again, the improvements in each metric confirm the advantage of CoPreFL.
@@
Table 5: Results with pre-training on a centrally stored public dataset. ImageNet is used for pre-training, while CIFAR-100 is used for downstream FL.
Application to centrally stored public dataset. We further consider the applicability of CoPreFL when the pre-training dataset is centrally stored, as discussed in Remark 2. To achieve this, we performed pre-training using the ImageNet_1K dataset and use FedAvg with CIFAR-100 as the downstream task. We intentionally split the dataset according to scenario I to mimic the distributed nature of downstream FL. The results are shown in Table 5, where we compare against standard centralized pre-training. The superiority in average accuracy and balanced performance achieved by CoPreFL confirms that CoPreFL is advantageous even when a large public dataset is used for pre-training. Detailed implementations and additional results can be found in Appendix D.7.
@@
Table 6: Results in the domain shift scenario. PACS dataset is used to model data discrepancy between pre-training and downstream stages.
Comparison under domain shifts. Our experiments so far have focused on the robustness of the pre-trained model to unseen labels. We now explore CoPreFL’s capability to handle unseen dataset domains during downstream tasks, using the PACS dataset. In Table 6, we use 3 domains (Art, Cartoon, and Photo) for pre-training in scenario I and conduct downstream FedAvg using the remaining Sketch domain, distributing samples across clients as in Table 1. These results show that CoPreFL’s improvements extend to cases when facing arbitrary downstream FL tasks with domain distributions that are distinct from those in the pre-training phase. This highlights CoPreFL’s ability to provide robust initializations when confronted with several different types of downstream data heterogeneity. We provide more detailed settings and results in Appendix D.1.

### v4

@@ 3.2 CoPreFL in Scenario I (Pre-training with Distributed Clients)
Model-agnostic meta update. Considering the objective function in (4), each participant $j$ downloads the temporary global model $\overline{\Phi^{t}}$ and employs its query set $Q_{j}$ to compute its local query loss $\mathcal{L}_{Q_{j}}(\overline{\Phi^{t}})$, as in line 15 in Algorithm 1. The gradients are also computed locally and sent back to the server, as both are necessary to conduct the meta-update. On the server-side, the overall query meta-loss $\mathcal{L}_{Q}(\overline{\Phi^{t}})$ and the performance variance $\sigma_{Q}^{2}(\overline{\Phi^{t}})$ are computed, according to (3). Then, CoPreFL updates the temporary pre-training model $\overline{\Phi^{t}}$ through a gradient step with the customized query meta-loss $\mathcal{L}_{meta}$ and the aggregated received gradients, to align it with (4). To derive the meta-loss $\nabla_{\Phi^{t-1}}\mathcal{L}_{meta}(\overline{\Phi^{t}})$, we express it through the chain rule as $\nabla_{\overline{\Phi^{t}}}\mathcal{L}_{meta}(\overline{\Phi^{t}})\times\frac{\partial\overline{\Phi^{t}}}{\partial\Phi^{t-1}}$. Writing $\overline{\Phi^{t}}=\sum_{j\in m}\frac{|S_{j}|}{\sum_{i\in m}|S_{i}|}\Phi^{t,E}_{j}=\sum_{j\in m}\frac{|S_{j}|}{\sum_{i\in m}|S_{i}|}(\Phi^{t,E-1}-\eta\nabla\ell^{E}_{S_{j}}(\Phi^{t}))$, it follows that
@@
$\begin{split}\nabla_{\Phi^{t-1}}&\mathcal{L}_{meta}(\overline{\Phi^{t}})=\nabla_{\overline{\Phi^{t}}}\mathcal{L}_{meta}(\overline{\Phi^{t}})\times\\ &\Big{(}1-\eta\sum_{j\in m}\frac{|S_{j}|}{\sum_{i\in m}|S_{i}|}\frac{\partial}{\partial\Phi^{t-1}}\nabla\ell^{E}_{S_{j}}(\Phi^{t})\Big{)}.\end{split}$ (5)
@@
If we ignore the second derivative term, the meta-loss gradient can be approximated as $\nabla_{\overline{\Phi^{t}}}\mathcal{L}_{meta}(\overline{\Phi^{t}})$. This is similar to making a first-order approximation to a meta-update, a common practice in the implementation of MAML variants to reduce complexity (Finn, Abbeel, and Levine 2017).
@@
The server then sends the meta-updated global model $\Phi^{t}$ to a new set of participants to begin the next round of pre-training. After $T$ rounds, the final global model $\Phi^{T}$ serves as the pre-trained model for initializing FL in the downstream tasks, i.e., in Figure 1, clients in any downstream task conduct FL starting from the pre-trained model $w^{0}=\Phi^{T}$.
@@
Remark 1 (Key characteristics of CoPreFL meta-update). Using the query datasets, CoPreFL applies a meta-update to the temporary pre-training model–a global model developed through FL on the support datasets. This differs significantly from existing meta-learning based FL methods, which update client models for personalization, as discussed in Section 2. Our method aims to tailor the pre-training model to adapt to any downstream FL tasks, addressing robustness against unseen and heterogeneous data, unlike existing personalization methods. As we will see in Section 4, this leads to notable improvements of CoPreFL over employing these prior methods for pre-training.
@@ 3.3 CoPreFL in Scenario II (Hybrid Client-Server Pre-Training)
We next explore a pre-training scenario where the server holds a small dataset $D^{s}$ drawn from the broader population distribution, alongside client-held data. Unlike scenario I, where client data was split into support and query sets, in scenario II, all client samples are used as support data, while the server’s data serves as the query set.
The procedure of CoPreFL for scenario II is detailed in Algorithm 2 in Appendix B. Here, we highlight the key differences from Algorithm 1. First, the temporary global model $\overline{\Phi^{t}}$ is aggregated from local models, each trained on a participant’s full local dataset $D^{p}_{j}$. Second, the meta-update of the temporary global model $\overline{\Phi^{t}}$ utilizes the server’s data. To help mimic the distributed nature of downstream FL tasks, we randomly split the dataset $D^{s}$ into $|m|$ equally to derive the average loss and variance objectives similar to (3). The temporary global model $\overline{\Phi^{t}}$ is then updated based on meta-loss $\mathcal{L}_{meta}(\overline{\Phi^{t}})$, calculated similarly to (4) but through meta-updates using the server’s partitioned data.
Note that unequal and/or non-uniformly random partitionings of the server-side dataset for meta-updating could be considered as alternatives. However, due to the server’s lack of prior knowledge about future downstream FL tasks during pre-training, including their dataset sizes and distributions, meta-updating the model with randomly allocated query sets is the most viable solution. We show in Section 4 that this partitioning provides significant performance improvements over other pre-training strategies.
@@
Remark 3.The theoretical link between pre-training strategies and downstream task performance remains an open problem. This challenge has existed both in centralized-to-centralized (Chang et al. 2020; Dong et al. 2023; Zhang et al. 2022; Hu et al. 2019; Yuan et al. 2024) and centralized-to-federated (Nguyen et al. 2023; Chen et al. 2023) transfers from pre-training to downstream, and persists in the distributed-to-federated case we consider here. We thus leave theoretical analysis of CoPreFL to future work, and instead validate its effectiveness through extensive experiments. The success of CoPreFL can largely be credited to advantage from meta-learning, which enhances robustness across diverse sets of downstream tasks.
@@ 4 Experiments
Table 1: Performance on 10 non-IID downstream FL tasks, initialized with various non-IID FL pre-training methods. Lowest $X$shows the average accuracy of clients with the lowest $X$accuracy. CoPreFL provides the best initialization across scenario, metric, and dataset.
@@ 4.1 Experimental Setup
Datasets and model. For evaluation, we use CIFAR-100 (Krizhevsky 2009), Tiny-ImageNet (Le and Yang 2015), FEMNIST (Caldas et al. 2018), and PACS (Li et al. 2017), following data splits provided in (Park et al. 2021), and adopt ResNet-18 (He et al. 2015). To model scenarios where downstream task labels are unknown during pre-training, we divide CIFAR-100 into 80 classes for pre-training and 20 for downstream tasks, and Tiny-ImageNet into 160 and 40 classes, respectively. We also explore mixed scenarios involving overlapping classes between pre-training and downstream tasks. Following (Yang, Chen, and Shen 2023; Zhang et al. 2020), we allocate 95% of the samples from the pre-training dataset for clients, and the remaining 5% form the server dataset. For PACS, we adopt a one-domain-leave-out setup (Li et al. 2022; Zhou et al. 2021) using different data domains for pre-training and downstream tasks. Detailed dataset information is available in Appendix C.1.
@@
Pre-training phase. We distribute the pre-training dataset to $|M|=100$ clients following non-IID data partitions according to a Dirichlet distribution (Morafah et al. 2022; Li, He, and Song 2021), and select $|m|=20$ participants out of the $|M|$ clients for each FL round. Results with different $|m|$ and IID setups are reported throughout Appendix D. We adopt a standard approach commonly used in meta-learning-based research (Jamal et al. 2020; Shu et al. 2019; Park et al. 2021) for support/query splitting, where we randomly partition each client’s data into 80% support and 20% query sets. See Appendix C.1-C.2 for more details.
@@
Downstream FL task and evaluation metrics. To generate each downstream FL task, we randomly select 5 of the 20 classes from the CIFAR-100 dataset and 40 classes from the Tiny-ImageNet dataset, and distribute the corresponding data samples to a set of $|G|=10$ clients following non-IID Dirichlet data distributions (see Appendix D for IID results). Each participant in the downstream phase utilizes 80% of its local data as training samples, while the remaining 20% is reserved for testing samples. We keep the training procedure consistent for each downstream task (see Appendix C.2 for detailed settings). For each task, we evaluate the final global model using test samples from each client $g\in G$, reporting the accuracy and variance of the accuracy distribution across the clients. We consider a total of $X=10$ downstream tasks, and the evaluation metrics are reported as averages (with standard deviations) across the tasks.
Data distribution. Data samples are distributed to $|M|=100$ clients for pre-training and $|G|=10$ clients for downstream FL tasks using the corresponding dataset based on a Dirichlet($\alpha$) distribution with $\alpha=0.5$, as done in the literature (Morafah et al. 2022; Li, He, and Song 2021).
@@
Baselines for pre-training. We compare CoPreFL with several established FL algorithms, including (i) standard FedAvg (McMahan et al. 2017), (ii) FedMeta (Chen et al. 2018), which employs meta-learning for unseen scenarios, and (iii) q-FFL(q $>$ 0) (Li et al. 2020), designed to balance performance across clients. When applying these baselines in scenario II, in each pre-training round $t$, after the global model $\Phi^{t}$ has been constructed, we further train with 5 additional iterations on the server dataset. This extended training follows the approach in (Yang, Chen, and Shen 2023; Bian et al. 2023), where the server’s data is used to further refine the global model. Similarly, we introduce a baseline called CoPreFL-SGD, which first constructs a global model according to CoPreFL and then further performs SGD iterations using server data on the global model. Finally, we also consider initializations based on (iv) conventional centralized pre-training (Nguyen et al. 2023), and popular FL algorithms like (v) SCAFFOLD (Karimireddy et al. 2019), (vi) FedDyn (Acar et al. 2021), and (vii) PerFedAvg (Fallah, Mokhtari, and Ozdaglar 2020) (see Table 3(a)).
Figure 2: Testing accuracy distributions in various non-IID FL tasks. CoPreFL achieves the best average accuracy (i.e., right-leaning distribution) and smaller performance variance (i.e., narrower distribution) while also improving worst-performing clients’ accuracies.
@@ 4.2 Experimental Results
Main results for scenarios I & II. Tables 1(a) and 1(b) present the test accuracies for CoPreFL across scenarios I and II on the CIFAR-100 and Tiny-ImageNet datasets. In scenario I, CoPreFL shows robust initializations for downstream FL tasks, obtaining a higher average accuracy and reduced performance variance across clients, and significantly improving accuracies for the worst-performing clients (Lowest 10-20% metrics). This highlights the benefits of balancing the competing objectives in (3) during meta-updates. See Appendix D.1 for supplemental analyses and results, including varying the number of participants in pre-training, and different data distributions in downstream FL tasks. In scenario II, CoPreFL also consistently outperforms baselines in this case by effectively utilizing server data in addition to balancing our objectives (3). The benefit of a small server-side dataset, when available, can be seen through the performance gains from Table 1(a) to 1(b). Further, the improvement over CoPreFL-SGD suggests that conducting SGD with server data in a centralized manner after meta-updating the global model may divert the pre-trained model away from our designed objectives. This emphasizes the importance of performing meta-learning on the server data following the partitioning outlined in Algorithm 2. For more results on various configurations, including impacts of varying server dataset sizes, see Appendix D.2.
Performance distribution comparison. Figure 2 show the testing accuracy distributions of the final global model on each client in downstream tasks. We visualize results for our method and those with the second-best average accuracy and the second-lowest variance from Table 1. CoPreFL shows consistently narrower distributions, reflecting reduced performance variance. Distribution shifts to the right can also be found, indicating higher average accuracy. Importantly, CoPreFL effectively shifts nearly all low-performing clients from the left tail of the baselines towards the right, enhancing accuracy for these clients. More results for various scenarios are available in Appendix D.3.
@@
Effect of balancer $\gamma$ in CoPreFL. Table 2 presents the performance of CoPreFL on Tiny-ImageNet using different balancers $\gamma$. A larger $\gamma$ implies that the pre-trained model prioritizes the devices’ average performance, whereas a smaller $\gamma$ emphasizes performance balance. We see that increasing $\gamma$ lead to higher average accuracy in downstream FL tasks but also greater variance, indicating performance imbalance. This trend shows that CoPreFL allows control over the relative importance between accuracy and balanced performance during pre-training.
@@
Centralized (Nguyen et al. 2023) 81.30 $\pm$ 2.92 69.44 $\pm$ 2.33
SCAFFOLD (Karimireddy et al. 2019) 79.15 $\pm$ 3.08 57.84 $\pm$ 1.95
FedDyn (Acar et al. 2021) 81.23 $\pm$ 2.96 53.17 $\pm$ 2.85
PerFedAvg (Fallah, Mokhtari, and Ozdaglar 2020) 81.58 $\pm$ 1.83 49.73 $\pm$ 2.65
@@
Comparison with other initialization methods. In addition to the baselines in Table 1(b) which consider performance balance or meta-learning, Table 3(a) also evaluates popular algorithms like SCAFFOLD (Karimireddy et al. 2019), FedDyn (Acar et al. 2021), and PerFedAvg (Fallah, Mokhtari, and Ozdaglar 2020) providing pre-training for scenario I. We see that CoPreFL also outperforms these baselines, validating our meta-learning approach based on (3). These popular FL algorithms struggle with unseen task heterogeneity and performance balance. We also explore initializing downstream FedAvg with random weights or a centrally pre-trained model, a concept introduced in (Nguyen et al. 2023). While the centralized method boosts downstream FL accuracy over random initialization, it introduces significant performance variance across clients, as it fails to mimic downstream FL characteristics. More details along with additional results can be found in Appendix D.5.
Compatibility with other downstream FL algorithms. We next explore the ability of our pre-training method to enhance the performance of downstream FL algorithms other than FedAvg. For this, we consider FedProx (Sahu et al. 2018) and q-FFL (Li et al. 2020), more advanced FL algorithms that addresses heterogeneity and performance balance, in each federated downstream task. Table 3(b) shows the results. Overall, we see that CoPreFL consistently achieves superiority in accuracy and variance compared to other pre-training baselines, when combined with different downstream FL algorithms. Details on the implementation and further discussions are provided in Appendix D.6.
@@
Table 4: Results with both seen/unseen classes during downstream FL, using the CIFAR-100 dataset in scenario I.
Both unseen/seen classes in downstream FL tasks. In addition to the setting without overlapping classes between pre-training and downstream tasks, we explore a mixed scenario where clients in the downstream FL also hold “seen classes.” To implement this, we use CIFAR-100 and randomly sampled 10 classes from the pre-training dataset and 10 classes from the downstream dataset, resulting in 10 seen and 10 unseen downstream classes. From this, we constructed 10 downstream tasks by randomly selecting 5 classes in each case, to conduct non-IID downstream FedAvg. The pre-trained models were solely trained on the original 80 classes of CIFAR-100. Table 4 shows the results. We see that the accuracies are generally higher compared to those in Table 1(a), as the downstream tasks involve classes seen during pre-training. Once again, the improvements in each metric confirm the advantage of CoPreFL.
@@
Table 5: Results with centrally stored dataset. ImageNet is used for pre-training, while CIFAR-100 is used for downstream FL.
Application to centrally stored public dataset. We also explore the applicability of CoPreFL with centrally stored pre-training data, as detailed in Remark 2. We conducted pre-training using the ImageNet_1K dataset and use FedAvg with CIFAR-100 as the downstream task. We intentionally split the dataset according to scenario I to mimic the distributed nature of downstream FL. Results in Table 5 show that CoPreFL outperforms standard centralized pre-training in both average accuracy and balanced performance, showing CoPreFL’s advantage even when a large public dataset is used for pre-training. Further details and additional results are available in Appendix D.7.
@@
Table 6: Results in the domain shift scenario using PACS dataset.
Comparison under domain shifts. Our experiments so far have focused on the robustness of the pre-trained model to unseen labels. We now explore CoPreFL’s ability to handle unseen data domains during downstream tasks using the PACS dataset. In Table 6, we use 3 domains (Art, Cartoon, and Photo) for pre-training in scenario I and conduct downstream FedAvg using the remaining Sketch domain, distributing samples across clients as in Table 1. Results show that CoPreFL effectively handles domain distributions in downstream FL tasks that differ from the pre-training phase. This highlights CoPreFL’s ability to provide robust initializations amid various types of downstream data heterogeneity. Additional settings and results are available in Appendix D.1.